OpenStudy (dc456):

can anyone help me with this please? Write the Riemann sum to find the area under the graph of the function f(x) = x4 from x = 5 to x = 7.

OpenStudy (solomonzelman):

\(x^4\) ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\limits_5^7x^4~dx }\) That is basically your question, and you would like to do this geometrically (right?). Are you given the method to use, and the number of shapes you must use, or anything?

OpenStudy (dc456):

that's all the question gives me

OpenStudy (solomonzelman):

So, you can do the right end, left end, sum, OR, upper sum, lower sum?? whichever one you choose?

OpenStudy (solomonzelman):

Well, I think that we can use n=4. \(\color{#000000 }{ \displaystyle \Delta x=\frac{b-a}{n}=\frac{7-5}{4}=\frac{1}{2} }\)

OpenStudy (dc456):

these are the choices

OpenStudy (solomonzelman):

Oh, there are options. This is different:)

OpenStudy (dc456):

Yes, should've added it with the question :P

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\limits_a^b f(x){\tiny~}{\mathrm d}x= \lim_{n\to\infty}\left[\sum_{i=0}^{n}f\left(x_i\right)\Delta x_i\right]}\) that is the general form; I think I got that right:)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \Delta x_i=\frac{a-b}{n}=\frac{7-5}{n}=\frac{2}{n}}\) \(\color{#000000 }{ \displaystyle c_i=b+\left(\frac{b-a}{n}\right)i=7+\left(\frac{7-5}{n}\right)i=\frac{2i}{n}}\)

OpenStudy (solomonzelman):

oh, that \(x_i\), I wrote it as \(c_i\).... I apologize for that. sorry

OpenStudy (dc456):

ok, I understand that part thank you

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle y(\omega) }\)

OpenStudy (solomonzelman):

Actually, to be precise and simple, you simply get: \(\color{#000000 }{ \displaystyle \int\limits_5^7{\tiny~}x^4{\tiny~}dx=\left.\frac{x^{4\color{red}{+1}}}{4\color{red}{+1}}\right|_{x{\tiny~~}={\tiny~~}5}^{x{\tiny~~}={\tiny~~}7} =\left.\frac{x^5}{5}\right|_{x{\tiny~~}={\tiny~~}\color{blue}{5}}^{x{\tiny~~}={\tiny~~}{\color{blue}{7}}} =\frac{(\color{blue}{7})^5}{5}-\frac{(\color{blue}{5})^5}{5}=2736.4}\)

OpenStudy (solomonzelman):

that is all you would normally need to do to find this area, using the \(\color{blue}{\rm power~~rule~~of~~integration}\), and by applying \(\color{blue}{\rm fundamental~~theorem~~of~~calculus}\).

OpenStudy (dc456):

thank you so much, I was so lost!

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\limits_a^b{\tiny~}x^n{\tiny~}dx=\left.\frac{x^{n\color{red}{+1}}}{n\color{red}{+1}}\right|_{x{\tiny~}={\tiny~}a}^{x{\tiny~}={\tiny~}b}=\frac{b^{n\color{red}{+1}}}{n\color{red}{+1}}-\frac{a^{n\color{red}{+1}}}{n\color{red}{+1}}}\) that is the power rule, by a definite integral.

OpenStudy (dc456):

would the answer be d?

OpenStudy (solomonzelman):

yes, for this purpose, yes.

OpenStudy (solomonzelman):

And the fundamental theorem of calculus \(\color{#000000 }{ \displaystyle \left. \int\limits_a^b{\tiny~}f'(x) {\tiny~}dx=f(x)\right|_{x~=~a}^{x~=~b}=f(b)-f(a)}\)

OpenStudy (dc456):

I understand, thank you! could I possibly ask you for help with one more?

OpenStudy (solomonzelman):

I will try, if I won't have to leave:)

OpenStudy (dc456):

thank you!

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \int\limits_0^{10}f(x)~dx }\) \(\color{#000000}{f(x)=\begin{cases} 5,\quad\quad\quad~ x\le 5 \\ 10-x,\quad x >5 \end{cases} }\)

OpenStudy (solomonzelman):

Can you please re-write this integral, as the sum of two functions, based on the piece wise function that you are given?

OpenStudy (solomonzelman):

(I mean as the sum of two integrals of two functions)

OpenStudy (dc456):

wouldn't they both equal 50?

OpenStudy (solomonzelman):

I don't know yet ... can you show me how you got this result please?

OpenStudy (dc456):

I did 5(10)-0

OpenStudy (solomonzelman):